OpenStudy (anonymous):

Help please! multiply (2 sqrt 5 - 1)(3 sqrt 5+4)

OpenStudy (anonymous):

\[Multiply. (2\sqrt{5}-1)(3\sqrt{5}+4)\]

OpenStudy (anonymous):

\[A) 26-11\sqrt{5}\] \[B) 26+5\sqrt{5}\] \[C) 34-11\sqrt{5}\] \[D) 34+5\sqrt{5}\]